OpenStudy (anonymous):

help me A real estate agent estimates that the value of office buildings in Lakewood will continue to go up by 5% per year. If an office building in this city is currently valued at $650,000, how much will it be worth in 10 years? If necessary, round your answer to the nearest cent.

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

@tom99

OpenStudy (133):

6500000

OpenStudy (anonymous):

a good equation is: \[Original Price \times Decimal Modifier ^{ Number of Changes}\] so in this case \[650000 \times 1.05^{10}\]

OpenStudy (anonymous):

650000 * 1.62889462678 = 1058781.50741 Then just round

OpenStudy (anonymous):

thank you so much ! @tom99

OpenStudy (anonymous):

Typewriters are going out of fashion in the modern era. If 343,000 typewriters were sold this year and sales decline at a rate of 15% per year, how many will be sold 14 years from now? If necessary, round your answer to the nearest whole number.

OpenStudy (anonymous):

@tom99